A medical call center has three types of callers (i.e. patients): the ones with a single request, the ones with multiple requests, and the ones with complex needs.

The processing times for the different types of callers (i.e. patients), in minutes, are listed in the following table.

Processing Time per patient (in minutes)

Type of patient

Demand (patients/day)

Receptionist

Nurse

Doctor

Single request

60

2

10

15

Multiple requests

40

4

20

20

Complex needs

30

4

0

30

There is one receptionist who answer all the calls, then directs then to a nurse, who then directs the patient call to a doctor. There are four nurses and five doctors. Everyone works 8 hours per day in this medical office.

What is the implied utilization (in percentage value, with no decimal -- for example, for 100%, type in 100) of the Nurses?
Answer #1

Total time in a day = 8*60 = 480 minutes

Total time nurse is occupied = 60*10+40*20+30*0 = 1400 minutes

Total nurses = 4

Implied Utilization = ( Occupied time)/ (# of resources * time in a day) = 1400/(4*480) = 72.92%